Assume that Ocean King Products sells three varieties of canned seafood with the following prices and costs.

Selling Price
per Case
Variable Cost
per Case
Fixed Cost
per Month
Variety 1 $ 16 $ 13
Variety 2 17 15
Variety 3 22 15
Entire firm $ 48,600

The sales mix (in cases) is 40 percent Variety 1, 35 percent Variety 2, and 25 percent Variety 3.

Required:

a. At what sales revenue per month does the company break even?

b. Suppose the company is subject to a 35 percent tax rate on income. At what sales revenue per month will the company earn $51,090 after taxes assuming the same sales mix?
Answer #1

Solution:

a)

Unit selling price of overall product E:

=(16 *0.4)+($17*0.35)+($22*0.25)

=6.4+5.95+5.5

=$17.85

Unit variable cost of overall product :

=(13*0.4)+(15*0.35)+(15*0.25)

=5.2+5.25+3.75

=14.2

Unit contribution margin E = Unit selling price E - unit variable cost of E

=$17.8 - $14.2

=$3.6

Break even point = fixed cost/contribution margin per unit

=$48,600 / 3.6

=$13,500

Break even point of variety 1= 13500*0.4= 5,400 cases

Break even point of variety 2= 13500*0.35= 4,725 cases

Break even point of variety 3= 13500*0.25= 3,375 cases

calculate sales revenue per month required to break even

Variety 1 sale revenue = 5,400 * $16 =$86,400

Variety 2 sale revenue = 4,725 * $17 =$80,325

Variety 3 sale revenue = 3,375 * $22 =$74,250

Total sales revenue = $86,400+$80,325+$74,250 =$240,975

b)

Target volume = fixed cost +[Target profit /(l-i)]/ contribution margin per unit

=$48,600+[$51,090/(1-.35)/]3.6

=$22,724 cases

Calculate the sales volume in units for each individuals product in their sales mix ratio

Target volume (cases) of variety1= 22,724 cases *0.4 = 9,089

Target volume (cases) of variety2= 22,724 cases *0.35 =7,953

Target volume (cases) of variety3= 22,724 cases *0.25 = 5,681

Calculate the sales revenue per month required by the company to ear an after tax operating profit of 51,090

Variety 1 target sale revenue = 9,089 *16 = 145,424

Variety 2 target sale revenue = 7,953 *17 = 135,201

Variety 3 target sale revenue = 5,681 *22 = 124,982

Total target sales revenue = $145,424+$135,201+$124,982 = 405,607